6. Problems and Applications Q6

Suppose Dmitri is the only seller in the market for bottled water and Antonio is the only buyer. The following lists show the value Antonio places on a bottle of water and the cost Dmitri incurs to produce each bottle of water:

Antonio's Value
Value of first bottle: $9
Value of second bottle: $7
Value of third bottle: $4
Value of fourth bottle: $1
Dmitri's Costs
Cost of first bottle: $1
Cost of second bottle: $4
Cost of third bottle: $7
Cost of fourth bottle: $9

The following table shows their respective supply and demand schedules:

Price

Quantity Demanded

Quantity Supplied

$1 or less 4 0
$1 to $4 3 1
$4 to $7 2 2
$7 to $9 1 3
More than $9 0 4

Use Dmitri's supply schedule and Antonio's demand schedule to find the quantity supplied and quantity demanded at prices of $2, $5, and $8. Enter these values in the following table.

Price

Quantity Demanded

Quantity Supplied

2
5
8

A price of   brings supply and demand into equilibrium.

At the equilibrium price, consumer surplus is

, producer surplus is

, and total surplus is

.

If Dmitri produced and Antonio consumed one less bottle of water, total surplus would   .

If instead, Dmitri produced and Antonio consumed one additional bottle of water, total surplus would   .
Answer #1

Here the quantity demanded and quantity supplied both are equal to 2 bottles at a price of $5 per bottles.

Consumer surplus = benefit received in terms of difference between valued money and money paid as price. At P = 5, this is (9 - 5) + (7 - 5) = $6

Producer surplus = benefit received in terms of difference between money received as price and valued money. At P = 5, this is (5 - 4) + (5 - 1) = $5

Total surplus = CS + PS = $11

If one less bottle was consumed, TS would have fallen from $11 to $8 ($9 - $5 + $5 - 1), a decline by $3.

If one more bottle was consumed, TS would have fallen from $11 to $6 ((9 - 5) + (7 - 5) + (4 - 5) + (5 - 7) + (5 - 4) + (5 - 1)), a decline by $2.

Price Quantity Demanded Quantity Supplied 2 3 1 2 2 5 1 3 8